# Poker Hand History File Format Specification

### v2

@@ VI-D Card Notation
TABLE V: All suits and their single character representations.
@@
Any card in the PHH format is either known or unknown. Each known card must be represented with two characters. The first character represents the rank of the card, and must either be one of the numeric characters from “2” to “9”, representing the analogous ranks from deuces to nines, or an uppercase alphabet character “T”, “J”, “Q”, “K”, and “A” for the ranks of tens, jacks, queens, kings, and aces, respectively. The second character represents the suit, the valid values of which must be the lowercase alphabet characters “c”, “d”, “h”, and “s” for clubs, diamonds, hearts, and spades, respectively. All ranks and suits are tabulated in Table IV and V, respectively. An unknown card must be represented with two question mark characters: “??”. Multiple cards must be represented by concatenating individual single-card representation without any separators or delimiters such as “AdAcKh” for ace of diamonds, ace of clubs, and king of hearts.
Certain variants like short-deck hold ’em use a non-standard deck where low ranks from deuces to fives do not exist. When a card that does not exist for a particular variant being played is specified, the parser should treat this as a violation of the specification.
@@ VII Optional Fields
TABLE VI: The optional fields and their TOML native types.
@@
Author author String
Event event String
URL url String
Address address String
City city String
Region region String
Postal code postal_code String
Country country String
Time time Local time
Time zone time_zone String
Day day Integer
Month month Integer
Year year Integer
Hand hand Integer
Level level Integer
Seats seats Array of integers
Number of seats seat_count Integer
Table table Integer
Players players Array of strings
Finishing stacks finishing_stacks Array of integers or floats
Currency currency String
Ante trimming status ante_trimming_status Boolean
Time limit time_limit Integer or float
Time banks time_banks Array of integers or floats
currency = # error
Figure 8: An example syntax error in PHH and TOML files [16].
@@
The optional fields may be left unspecified by the annotator. The complete list of optional fields and their TOML native types are tabulated in Table VI. When a field is included, it must be completely defined. For instance, a partial definition as in Figure 8 is strictly forbidden as it is syntactically incorrect as per the TOML specification [16].
VII-A Annotator
VII-A1 Author
The “author” field describes the annotator and is of TOML native string type. The value should represent a properly capitalized full name of the annotator as in
``$First\_name[\ Middle\_name(s)...]\ Last\_name"$
or a mononym
``$Name"$
depending on their preferences.
VII-B Venue
The following fields describe the venue in which the hand was played.
VII-B1 Event
The “event” field represents the full name of the event which should be in English and is of type: string.
VII-B2 URL
The “url” field contains a string URL relevant to the event. This field is particularly important for online poker hands where the venue might not be a physical location. Still, the annotator may choose to specify a relevant URL to the organizer’s website or tournament website for physical poker hands, as well. The link should be prefixed with substrings like “http://” or “https://” to denote the internet communication protocol.
VII-B3 Address
The “address” field contains a string value that may or may not include the street name, street number, room number, et cetera. The annotator should use English wherever possible.
VII-B4 City
The “city” field contains the information about the city the hand was played in. The annotator should use full English names for the cities.
VII-B5 Region
The “region” field contains the information about the city the hand was played in. Similarly to the cities, the annotator should use full English names for the regions.
VII-B6 Postal Code
The “postal_code” field contains the postal code of the location of the hand.
VII-B7 Country
The “country” field contains the name of the country of the hand location. Again, the annotator should use full English names for the countries.
VII-B8 Time
The “time” field should contain the approximate time the hand began as a TOML native local time without the timezone information.
If the “time_zone” field is defined, the time is assumed to be of that time zone. Otherwise, if every one of the following three fields – “city,” “region,” and “country” – are defined, it is assumed that the time is in the local time zone of the location of the hand. If the two previous conditions are not satisfied, the time is assumed to be in Coordinated Universal Time (UTC).
VII-B9 Time Zone
The “time_zone” field contains a value of type string and must represent an IANA time zone in the latest version of the IANA time zone database [17] available like “America/Toronto.”
VII-B10 Day
The “day” field should contain the day the hand took place in, and is of the integer type.
VII-B11 Month
The “month” field should contain the month the hand took place in, and has an integral value where 1 represents January, 2 represents February, and so on.
VII-B12 Year
The “year” field should contain the year the hand took place in, and must be an integer.
VII-C Counter
VII-C1 Hand
The “hand” field denotes the hand number which should be an integer.
VII-C2 Level
The “level” field denotes the blind or ante level, which is relevant in a poker tournament. This field must contain an integer value.
VII-D Seating
VII-D1 Seats
The “seats” field denotes the seat numbers of the players. The value of this field must be an array of integers, of length $n$ where $n$ is the number of players.
VII-D2 Number of Seats
The “seat_count” field denotes the number of seats and must have an integer value. Note that it is possible for there to be more seats than the players.
VII-D3 Table
The “table” field denotes the table number the hand is played in, as an integer value.
VII-E Miscellaneous
VII-E1 Players
The “players” field contains player names as an array of strings. The names should be written as a full name (“First_name[ Middle_name(s)…] Last_name”) or as a mononym (“Name”) depending on which is used or preferable. If the name of the player is unavailable perhaps due to anonymity, an empty string should be used.
VII-E2 Finishing Stacks
The “finishing_stacks” field denotes the final stacks as an array of non-negative integers or floats. It may also be specified for a non-terminal hand history file, which can be interpreted as the stack values after all the action notations are applied.
@@ VII-E2 Finishing Stacks
The description of finishing stacks is helpful as the parser may not be aware of the granularity of the currency the chips are in. For example, dollars can be broken down to cents whereas Japanese Yen must be of an integral value. On top of this, in a physical setting where chips are used, depending on the denominations, odd chip situations may arise where the player out of position is given the extra odd chip that cannot be broken further. It is, of course, infeasible to describe all the different chip values each player has in a poker hand history format. These inaccuracies are inherent drawbacks of using purely numerical representations to describe the stack values. It is worth noting that the inconsistencies caused by such circumstances only lead to extremely minor ambiguities in the final stack sizes that should not significantly impact the expected value calculations.
VII-E3 Currency
The “currency” field denotes what currency the chips are in. The value must be of string and be one of three letter currency values in ISO 4127 standard [18].
VII-E4 Ante Trimming Status
The “ante_trimming_status” denotes how to handle the special cases where a player or players are so short-stacked that they cannot even pay the full ante as a Boolean value that defaults to “false.” If “true,” the player can only win depending on how much the player contributed. Otherwise, the player can win all the antes even if they did not pay the full ante. This field was introduced due to the ambiguities in the WSOP rules [6, 7]. Note that, in certain ante configurations such as big blind antes, unequal contribution in the antes is expected, and therefore this value should be kept as “false” for non-uniform antes. In the vast majority of real-life poker hands, the players can at least pay the full antes. Therefore, this field is only relevant in extremely rare circumstances.
VII-F Time Control
VII-F1 Time Limit
The “time_limit” represents the shot clock time and is of type: integer or float. It represents the time the user has to make a decision at every turn. If there is none, the annotator can omit this field or set it as “inf” for infinity.
VII-F2 Time Banks
The “time_banks” field denotes the time banks of each player at the beginning of the hand, as an array of integers or floats of length equal to the number of players. Just like in stack values, representing time banks as a simple numerical value loses the detail of the granularity of the time cards a player may have. Still, integral values are used thanks to their simplicity and consistency.
@@ VIII User-defined Fields
_sponsors = ["", "GGPoker", ""]
_nationalities = [
  "Peru", "Italy", "Japan",
Figure 9: Example user-defined fields in PHH format.
Depending on the use case, the required and optional fields may not be sufficient to capture certain desired levels of details, as demonstrated in Figure 9. The specification allows annotators to define custom fields. Parser implementations should simply ignore unknown fields and may issue a warning when such a field is encountered. This is also in part to ensure that the parser is compatible with PHH files conforming to the future version of the specification that may introduce new fields.
The names of the user-defined fields must not be one of the names of the required or optional fields and should be chosen carefully such that it does not conflict with future versions of the specification. The naming conventions of user-defined fields are explained in the style guide.
IX Style Guide
In PHH files, the fields may be in any order, and the formatting may follow any style, as long as it conforms to the TOML syntax [16]. However, the annotator is encouraged to list all required fields first before listing the optional fields and enclose the equal signs “=” for each key/value assignment with a single space character “ ” on both sides to enhance readability.
When naming the user-defined fields, users should use the “snake_case” naming convention where words are written in lowercase characters, separated by an underscore as in the required and optional fields to maintain consistency. In addition, the names of the user-defined fields should start with a single underscore as shown in Figure 9 to avoid conflicts with the future version of the specification.
The PHH files should have a “.phh” filename extension. As for the file naming conventions of standalone hands, we suggest including the last names of the relevant involved players in lowercase characters, with the last names sorted in alphabetical order, each separated by a single dash “-” character, followed by the year the hand took place. For example, the file containing the hand in Figure 1 can be named “dwan-ivey-2009.phh” whereas “arieh-yockey-2019.phh” can be the name of the PHH hand file in Figure 2. As for a bulk of hands for a particular event, the PHH files should all be placed in a folder or nested folders. For an event like 2023 WSOP Event #43: \$50,000 PPC – Day 5 [11], we suggest “wsop/2023/43/5” as the nested folders’ name. For filenames of the files in such folders, we suggest simply using its hand number like “10.phh” for Hand 10 when there is only one table (final table in this case), or using both the hand number and the table number as in “6-51.phh” for Hand 51 in Table 6. These naming conventions are followed in the supplementary PHH files.
X Benchmarking
TABLE VII: The conciseness of each hand history format.
@@ X Benchmarking
Using the average number of newlines, words, and bytes used to represent a hand, we compare the conciseness of the following three formats: the PHH format, the PokerStars hand history format, and the format used in the supplementary of Brown and Sandholm [5] in Table VII. The dataset of hands explored in this analysis is the 10,000 hands played by the poker AI agent Pluribus, all of which were of the no-limit Texas hold ’em variant. The PokerStars hand history version of the dataset was provided by Wang [19] and is used for this comparison. All files exclusively use the ASCII character set.
@@
In order to validate the correctness of the PHH format and our open-source parser implementation, the parsed games were checked for consistency and correctness by comparing the final stacks with the payoffs provided in the dataset of Brown and Sandholm [5].
XI Tooling
from pokerkit import *
# Hand loading
with open("...", "rb") as file:
    hh = HandHistory.load(file)
# Iterate through each action step
for state in hh:
Figure 10: Loading a PHH file in PokerKit [8].
from pokerkit import *
A = Automation
# Game state construction
game = PotLimitOmahaHoldem(
        A.ANTE_POSTING,
        A.BET_COLLECTION,
        A.BLIND_OR_STRADDLE_POSTING,
        A.CARD_BURNING,
        A.HOLE_CARDS_SHOWING_OR_MUCKING,
        A.HAND_KILLING,
        A.CHIPS_PUSHING,
        A.CHIPS_PULLING,
    True,
    0,
    (500, 1000),
    1000,
state = game((1259450.25, 678473.5), 2)
# State progression; Pre-flop
state.deal_hole("Ah3sKsKh")  # Antonius
state.deal_hole("6d9s7d8h")  # Blom
# Blom
state.complete_bet_or_raise_to(3000)
# Antonius
state.complete_bet_or_raise_to(9000)
# Blom
state.complete_bet_or_raise_to(27000)
# Antonius
state.complete_bet_or_raise_to(81000)
state.check_or_call()  # Blom
# Flop
state.deal_board("4s5c2h")
# Antonius
state.complete_bet_or_raise_to(91000)
# Blom
state.complete_bet_or_raise_to(435000)
# Antonius
state.complete_bet_or_raise_to(779000)
state.check_or_call()  # Blom
# Turn & River
state.deal_board("5h")
state.deal_board("9c")
# Creating hand history
hh = HandHistory.from_game_state(
    game, state,
hh.players = [
    "Patrik Antonius", "Viktor Blom",
# Dump hand
with open("...", "wb") as file:
    hh.dump(file)
Figure 11: Dumping to a PHH file in PokerKit [8]. This hand, played between Viktor Blom and Patrik Antonius, holds the record for the largest online poker pot ever played [20].
From PokerKit [8] Version 0.4, the library features PHH file format reader and writer utilities. It offers “load” and “dump” programmatic APIs akin to those provided by Python’s standard libraries such as “json,” and “pickle” [21]. Figures 10 and 11 show sample usages of the PHH file format utilities in PokerKit. The hand history object in Python serves as an iterator of the corresponding poker state which first yields the initial state, followed by the same state after applying each action one-by-one in the “actions” field. From game and state objects that are interacted with programmatically, the hand history object can also be created which can subsequently be saved in the file system.
XII Discussion
@@
XII-A AI Agents
The use of the PHH format can greatly benefit AI research, especially those that use reinforcement learning and ensemble learning techniques. The comprehensive hand histories allow for detailed analyses and can be used for training purposes. Furthermore, using a standardized format to share hand histories played by AI agents can also enhance transparency and enable other researchers to easily validate their findings.
XII-B Online Casinos
Online casinos that offer poker games can store hands played on their platforms in the PHH file format for archival purposes that can track player history and facilitate easier internal auditing processes. The same data (or maybe another version with the sensitive information like other players’ hole cards censored out) can be provided to the players involved in the hand upon request. As an alternative to using persistent files, the developers can map each field into database table columns to store them in a queryable database.
XII-C Poker Tools
@@
XIII Conclusion
This paper introduces the PHH file format, designed to concisely describe poker hand history while allowing ease in human usage and machine parser implementations. Our solution fills a long-standing gap in the existing literature on the computer poker field. The specification builds on top of the TOML specification [16], allowing software systems to leverage the rich ecosystem surrounding the TOML file format. By representing different game state information through fields, the hand can be described in a consistent, structured manner. We provide a demonstration of our open-source Python PHH hand parser implementation that utilizes the PokerKit [8] framework as a hand simulator. In the supplementary, we provide over 10,000 PHH files in 11 different diverse variants to validate the format’s flexibility and effectiveness.
The PHH format is suitable to be used in a variety of poker applications in use cases ranging from poker AI development, hand archival, analytical tools, and teaching materials. A promising future direction for research in the field of computational poker is in the development of a poker engine communication protocol akin to the Universal Chess Interface for chess [22] which may utilize numerous portions of the PHH format. The development of such a standard has the potential to accelerate the development of industrial poker AI tools and facilitate their direct comparisons.

### v3

@@ VI-A Card Notation
TABLE VI: All suits and their single character representations.
@@
A card in the PHH format can either be known or unknown. Each known card must be represented with two characters. The first character represents the rank of the card while the second character represents the suit, as tabulated in Tables V and VI. An unknown card must be represented with two question mark characters: “??”. Multiple cards must be represented by concatenating individual single-card representations without any separators or delimiters.
@@ VII Optional Fields
TABLE VII: The optional fields and their TOML native types.
@@
Annotator full name or mononym author String
Event name event String
Event or organizer URL url String
Venue street-level address address String
Venue city city String
Venue region, state, or province region String
Venue postal code postal_code String
Venue country country String
Timestamp at the start of the hand time Local time
IANA time zone time_zone String
Event day day Integer
Event month month Integer
Event year year Integer
Hand number hand Integer
Tournament level level Integer
Players’ seat numbers seats Array of integers
The number of seats seat_count Integer
Table number table Integer
Player full names or mononyms players Array of strings
Final stacks finishing_stacks Array of integers or floats
ISO 4127 [18] currency currency String
Ante uniformity ante_trimming_status Boolean
Allocated time per action time_limit Integer or float
Time banks at the beginning of the hand time_banks Array of integers or floats
@@
The optional fields may be left unspecified by the annotator. The complete list of optional fields and their TOML native types are tabulated in Table VII.
@@
The description of finishing stacks is helpful as the parser may not be aware of the granularity of the currency the chips are in or the rake and rake-cap applied in the end. On top of this, in a physical setting where chips are used, depending on the denominations, odd chip situations may arise where the player out of position is given the extra odd chip that cannot be broken further. It is, of course, infeasible to describe all the different chip values each player has in a poker hand history format. These inaccuracies are inherent drawbacks of using purely numerical representations to describe the stack values. It is worth noting that the inconsistencies caused by such circumstances only lead to extremely minor ambiguities in the final stack sizes that should not significantly impact the expected value calculations.
@@ VIII User-defined Fields
Depending on the use case, the required and optional fields may not be sufficient to capture certain desired levels of detail. The specification allows annotators to define custom fields with a single underscore prefix in the field names. This is also in part to ensure that the parser is compatible with PHH files conforming to the future version of the specification that may introduce new fields.
IX Benchmarking
TABLE VIII: The conciseness of each hand history format.
@@ IX Benchmarking
Using the average number of newlines, words, and bytes used to represent a hand, we compare the conciseness of the following three formats: the PHH format, the PokerStars hand history format, and the format used in the supplementary of Brown and Sandholm [5] in Table VIII. The dataset of hands explored in this analysis is the 10,000 hands played by the poker AI agent Pluribus, all of which were of the no-limit Texas hold ’em variant. The PokerStars hand history version of the dataset was provided by Wang [19] and is used for this comparison. All files exclusively use the ASCII character set.
@@
Note that the benchmarks of this sort for other hands included in the supplementary is impossible due to the fact that the Pluribus format only supports no-limit Texas hold ’em in specific initial configurations and no documentation is available on recording hands of different variants and configurations in PokerStars notation. In addition, the concepts used in some of the histories like big blind antes are only recently emerging and therefore not supported on the PokerStars platform.
TABLE IX: The performance of the open-source PHH file format parser.
Format Throughput (hands/s)
PHH 6801.31
The performance of our open-source parser implementation, written in Python, is tabulated in Table IX. The hands in the supplementary were used to calculate the throughput and the parser process was run on a single CPU of Intel® Core™ i5-4690. All the file contents were loaded into memory prior to being parsed. We utilize Python’s built-in TOML file reader to parse the file and perform checks to validate the data. On average, it takes 0.147 milliseconds to parse a single hand. Note that no widely established open-source hand history parser for other file formats is available to facilitate a fair comparison.
X Validation
We performed multiple experiments in order to validate the correctness of the PHH format, the sample hand histories, and our open-source parser implementation.
First, all the hands included in the supplementary were validated through round-trip testing where the files were converted into snapshots of game states at each action step. These snapshots were then converted back into hand files and compared against the originals. Next, we developed a poker bot evaluation platform with 4 fixed-policy agents biased toward different poker actions and an agent connected to the Slumbot [20] API, the winning heads-up poker agent in the 2012 Annual Computer Poker Competition. Overall, these agents competed against each other for over ten thousand hands. These hands were then saved and the same round-trip testing method was carried out to these files.
Additionally, the parsed games for the Pluribus [5] hands were checked for consistency and correctness by comparing the final stacks with the payoffs provided in the dataset of Brown and Sandholm [5]. Similarly, the most of the televised poker hands from our historical selections and the final table of the 2019 WSOP Event #52: \$50,000 PPC on Day 5 had overlays of final player stacks in the source video from which we verified that the actions lead to the corresponding stack values. The actions were also checked for violations of the 2023 WSOP Tournament Rules [6].
XI Discussion
@@
XI-A AI Agents
The use of the PHH format can greatly benefit AI research, especially those that use reinforcement learning and ensemble learning techniques. The comprehensive hand histories allow for detailed analyses and can be used for training purposes. As previously mentioned, previous AI agents were greatly limited in scope due to the initial state assumptions they made about the game. Development of this file format serves as the crucial first step to set up the development of the poker AI agents that overcome these limitations. Furthermore, using a standardized format to share hand histories played by AI agents can also enhance transparency and enable other researchers to easily validate their findings.
XI-B Online Casinos
Online casinos that offer poker games can store hands played on their platforms in the PHH file format for archival purposes that can track player history and facilitate easier internal auditing processes. In the industry, logs of the game actions or table data at each snapshot are typically stored. This makes it difficult to carry out statistical analyses or automate certain tasks that may be of interest to these firms. An area of particular interest in online poker is cheating detection. A more highly structured data format for hand histories can facilitate the development of new tools or methods to discern unusual behaviors and identify tell-tale signs of bots. The same data (or maybe another version with the sensitive information like other players’ hole cards censored out) can be provided to the players involved in the hand upon request. As an alternative to using persistent files, the developers can map each field into database table columns to store them in a queryable database.
XI-C Poker Tools
@@
Crowdsourcing and opening of such a database to the public can open up new avenues of poker strategy research. There are largely two schools of poker playing (and developing imperfect information game AI agents): exploitative and unexploitative plays. The first school – the exploitative approach – advocates for modeling opponents’ behaviors and adjusting one’s strategy to take advantage of imbalance in the opponent’s range. The second school advocates for adopting a perfectly balanced strategy that cannot be exploited – Nash Equilibrium. Professional players generally advocate for a mix of both, while academics have mostly focused on the second school of calculating Nash equilibrium strategies [1, 2, 3, 4, 5]. Commonly mentioned issues with resolving of Nash equilibrium are that there may exist multiple Nash equilibriums and that it is not well defined in games of more than 2 players due to possibilities of collusion or cooperation among players. A large database of hand histories can provide data to model an average poker player from which exploitative AI agents can be developed with a more well-defined goal of taking maximum advantage of imbalances in commonly employed human poker strategies by diverging from Nash equilibrium.
XII Conclusion
This paper introduces the PHH file format, designed to concisely describe poker hand history while allowing ease in human usage and machine parser implementations. Our solution fills a long-standing gap in the existing literature on the computer poker field. The specification builds on top of the TOML specification [17], allowing software systems to leverage the rich ecosystem surrounding the TOML file format. By representing different game state information through fields, the hand can be described in a consistent, structured manner. In the supplementary, we provide 10,088 PHH files in 11 different diverse variants to validate the format’s flexibility and effectiveness.
The PHH format is suitable to be used in a variety of poker applications in use cases ranging from poker AI development, hand archival, analytical tools, and teaching materials. A promising future direction for research in the field of computational poker is in the development of automated systems that can digitalize televised poker hands into a formats, such as the PHH format, for further analysis and study. This technology would leverage computer vision and machine learning algorithms to recognize chip amounts, player actions, and dealt cards from video footages and/or onboard graphics. Such systems could greatly enhance the efficiency and accuracy of data collection for poker research, eliminating the manual entry of hand histories and reducing human error.